# An operator splitting analysis of Wasserstein–Fisher–Rao gradient flows

### v1

@@ 2.2 Variational Formula: FR-W
The reverse scheme described in (2.7) requires a more involved analysis than in Proposition 2.2, due to the lack of invertibility of $S_{\textrm{W}}$ in general. We instead rely on techniques used in developing Alekseev-Groebner type formulae, in particular, the formal Lie calculus, as is common in the analysis of operator splitting schemes [descombes_lietrotter_2013, Hundsdorfer2003]. Once again, we rely on the shorthand notation $\eta_{\gamma}$ to refer to $\eta_{1}(x;\gamma)$.
@@
Sequential split PDE: FR-W. Let us denote $\eta_{\gamma}=\eta_{1}(x;\gamma)$ and let Assumption 1 hold. The variation of one sequential split step of size $\gamma$ in the order FR-W (2.7) corresponds to the PDE
$\partial_{\gamma}\eta_{\gamma}=f_{\textrm{W}}(\eta_{\gamma})+f_{\textrm{FR}}(\eta_{\gamma})+\int_{0}^{\gamma}S_{\textrm{W}}\left(\gamma-\tau,-f_{\textrm{P}}(\eta_{\tau})\right)d\tau,$ (2.13)
@@
$\partial_{\gamma}\eta_{\gamma}=f_{\textrm{W}}(\eta_{\gamma})+f_{\textrm{FR}}(\eta_{\gamma})+\sum_{k=1}^{\infty}\frac{\gamma^{k}}{k!}\eta_{\gamma}\left(g_{k}(\eta_{\gamma})-\mathbb{E}_{\eta_{\gamma}}[g_{k}]\right),$ (2.14)
@@
$g_{0}(\eta_{\gamma})$ $=-g(\eta_{\gamma})=-\log\frac{\eta_{\gamma}}{\pi},$
$g_{k}(\eta_{\gamma})$ $=-g_{k-1}^{\prime}f_{\textrm{W}}(\eta_{\gamma})+\nabla\log\frac{\eta_{\gamma}}{\pi}\cdot\nabla g_{k-1}(\eta_{\gamma})+\frac{1}{\eta_{\gamma}}\nabla\cdot(\eta_{\gamma}\nabla g_{k-1}(\eta_{\gamma})),\quad\text{for}\enskip k=2,3,\dots$
and $g_{k-1}^{\prime}f_{\textrm{W}}(\eta_{\gamma})$ denotes the directional derivative of $g_{k-1}$ at $\eta_{\gamma}$ in the direction $f_{\textrm{W}}(\eta_{\gamma})$.
Proof.
By similar reasoning as in Section 2.1, the sequential split solution at time $\gamma$ in the order FR-W is given by $\eta_{\gamma}=S_{\textrm{W}}(\gamma,S_{\textrm{FR}}(\gamma,\mu_{0}))$ and its first variation w.r.t. $\gamma$ can be computed as
$\partial_{\gamma}\eta_{\gamma}=f_{\textrm{W}}(\eta_{\gamma})+\partial_{2}S_{\textrm{W}}(\gamma,u_{\gamma})f_{\textrm{FR}}(u_{\gamma}),$ (2.15)
where $u_{\gamma}:=S_{\textrm{FR}}(\gamma,\mu_{0})$. By Lemma 1 in [descombes_lietrotter_2013] and recalling that $S_{\textrm{W}}$ is a linear operator for which $\partial_{2}S_{\textrm{W}}(t,v)h=S_{\textrm{W}}(t,h)$, we have
$\partial_{2}S_{\textrm{W}}(\gamma,u_{\gamma})f_{\textrm{FR}}(u_{\gamma})$ $=f_{\textrm{FR}}(S_{\textrm{W}}(\gamma,u_{\gamma}))+\int_{0}^{\gamma}\partial_{2}S_{\textrm{W}}(\gamma-\tau,u_{\gamma})[f_{\textrm{W}},f_{\textrm{FR}}](S_{\textrm{W}}(\tau,u_{\gamma}))d\tau$
$=f_{\textrm{FR}}(\eta_{\gamma})+\int_{0}^{\gamma}S_{\textrm{W}}\left(\gamma-\tau,[f_{\textrm{W}},f_{\textrm{FR}}](S_{\textrm{W}}(\tau,u_{\gamma}))\right)d\tau,$ (2.16)
since $S_{\textrm{W}}$ is a linear operator and where $f_{\textrm{W}},f_{\textrm{FR}}](v)$ refers to the Lie commutator (see (3.2a) in [descombes_lietrotter_2013]),
where $D_{\textrm{FR}}$ is the Lie derivative operator associated with $f_{\textrm{FR}}$, i.e. $D_{\textrm{FR}}H(v)=H^{\prime}(v)f_{\textrm{FR}}(v)$ for any unbounded nonlinear operator $H:X\rightarrow\mathcal{T}$, $X\subset\mathcal{P}_{2}^{ac}$ and $H^{\prime}(v)$ denotes the Fréchet derivative operator evaluated at $v$. We note that while $f_{\textrm{W}}$ and $f_{\textrm{FR}}$ may not be Fréchet differentiable at any arbitrary $v$, their directional derivative exists and is sufficient to define the Lie derivative, therefore we work with the directional derivative. With a slight abuse of notation, we use $f_{\textrm{W}}^{\prime}f_{\textrm{FR}}(v)$ in place of $f_{\textrm{W}}^{\prime}(v)f_{\textrm{FR}}(v)$. Substituting (2.16) into (2.15) yields that the variation of $\eta_{\gamma}$ takes the form of the usual WFR PDE plus a perturbation term,
$\partial_{\gamma}\eta_{\gamma}=f_{\textrm{W}}(\eta_{\gamma})+f_{\textrm{FR}}(\eta_{\gamma})+\int_{0}^{\gamma}S_{\textrm{W}}\left(\gamma-\tau,[f_{\textrm{W}},f_{\textrm{FR}}](S_{\textrm{W}}(\tau,u_{\gamma}))\right)d\tau,$ (2.19)
where $f_{\textrm{W}},f_{\textrm{FR}}]S_{\textrm{W}}(\tau,u_{\gamma})\in\mathcal{T}$. Substituting in (A.6) developed in Lemma A.2 for the Lie commutator leads to (2.13).
$\int_{0}^{\gamma}S_{\textrm{W}}\left(\gamma-\tau,[f_{\textrm{W}},f_{\textrm{FR}}](S_{\textrm{W}}(\tau,u_{\gamma}))\right)d\tau=\sum_{k=1}^{\infty}\frac{\gamma^{k}}{k!}\mathcal{G}^{k}(\eta_{\gamma}).$
The result then follows using Lemma A.2. ∎
The expression in (2.14) is obtained by writing the perturbation term in (2.13) using the Lie commutator in (2.17) and higher order commutators. We work with this formulation as the inverse of the solution operator does not lead to a tractable form, when it exists. A similar approach can be used to obtain the perturbation term in Proposition 2.2 with $g_{k}=|\nabla g|^{2}$ for all $k=1,2,3,\dots$. We do not use this approach in the proof of Proposition 2.2 but exploit the invertibility of the unnormalised FR solution operator.
As Proposition 2.3 shows, the evolution of the splitting solution obtained with FR-W does not correspond to that obtained in Proposition 2.2 for W-FR. In the multivariate Gaussian setting, i.e. with $\mu_{0}(x)=\mathcal{N}(x;m_{0},C_{0})$ and $\pi(x)=\mathcal{N}(x;m_{\pi},C_{\pi})$, a more interpretable PDE reminiscent of (2.9) can be obtained (see Appendix B.2 for a derivation),
$\partial_{\gamma}\eta_{\gamma}$ $=f_{\textrm{W}}(\eta_{\gamma})+f_{\textrm{FR}}(\eta_{\gamma})-\eta_{\gamma}\left(g_{\text{FRW}}(\eta_{\gamma})-\mathbb{E}_{\eta_{\gamma}}[g_{\text{FRW}}(\eta)]\right),$ (2.20)
@@ 3 Splitting & convergence speed: multivariate Gaussian case
where $\Gamma:=C_{\pi}^{-1}+\frac{1}{2}I$. A derivation of (3.1)-(3.2) can be found in Lemma B.2 (letting $M_{t}=0$), which extends the derivation of $C_{t}^{-1}$ provided in [liero2025evolution]. To the best of our knowledge, the analytic solution for $m_{t}$ provided in (3.2) has not yet appeared in the literature.
Similarly, the splitting schemes yield Gaussian solutions at each time, as it is well-known that both W and FR flows preserve Gaussianity. We denote by $b_{n\gamma},Q_{n\gamma}$ the mean and covariance respectively at step $n$ of the W-FR split scheme with step size $\gamma$ and $a_{n\gamma},P_{n\gamma}$ the mean and covariance at step $n$ of the FR-W split. We write $b_{\gamma},Q_{\gamma}$ and $a_{\gamma},P_{\gamma}$ when $n=1$.
The similarity in structure of the sequential split PDEs (2.9) and (2.20) in the Gaussian setting also allows for obtaining analytic solutions of the mean and covariance in a similar fashion. By an application of Lemma B.2 with $\gamma$ in place of $t$ and setting $M_{\gamma}=(e^{\gamma}-1)I$, one obtains
$Q_{\gamma}$ $=C_{\pi}+(e^{\Gamma\gamma}(C_{0}-C_{\pi})^{-1}e^{\Gamma\gamma}+(e^{\gamma}-1)C_{\pi}^{-1})^{-1},$ (3.3)
$b_{\gamma}$ $=m_{\pi}+(Q_{\gamma}-C_{\pi})e^{\gamma C_{\pi}^{-1}}(C_{0}-C_{\pi})^{-1}(m_{0}-m_{\pi}).$ (3.4)
for the mean $b_{\gamma}$ and covariance $Q_{\gamma}$ of a single step of W-FR of size $\gamma$. Similarly, setting $M_{\gamma}=-\frac{1}{2}C_{\pi}(e^{2\gamma C_{\pi}^{-1}}-I)$ in Lemma B.2 yields
$P_{\gamma}$ $=C_{\pi}+\left(e^{\Gamma\gamma}(C_{0}-C_{\pi})^{-1}e^{\Gamma\gamma}+(e^{\gamma}-1)C_{\pi}^{-1}e^{2\gamma C_{\pi}^{-1}}\right)^{-1}$ (3.5)
$a_{\gamma}$ $=m_{\pi}+(P_{\gamma}-C_{\pi})e^{\gamma C_{\pi}^{-1}}(C_{0}-C_{\pi})^{-1}(m_{0}-m_{\pi}).$ (3.6)
for the mean $a_{\gamma}$ and covariance $P_{\gamma}$ of a single step of FR-W of size $\gamma$.
The structure of the mean for both splitting schemes is exactly the same as that obtained for the exact WFR flow with $C_{t}$ replaced by $Q_{\gamma},P_{\gamma}$ respectively. This demonstrates that in the Gaussian case the speed-up due to splitting is driven primarily through the improvement in covariance estimation. In particular, notice the difference in structure between $C_{t}$ vs $Q_{\gamma}$ and $P_{\gamma}$. The difference due to operator ordering is more subtle, as $Q_{\gamma}$ and $P_{\gamma}$ differ only by a single factor $e^{2\gamma C_{\pi}^{-1}}$. As will be shown later in this section, this factor does not necessarily lead to acceleration in all cases, and depends on the relationship between $\mu_{0}$ and $\pi$.
@@
Before presenting the main analytic results, consider Figure 3.1 which compares KL distance from $\pi$ after a single sequential split step and the same for the exact WFR in the univariate Gaussian setting. More specifically, it shows $\mathrm{KL}(\nu_{\gamma}||\pi)-\mathrm{KL}(\mu_{\gamma}||\pi)$ (blue line) and $\mathrm{KL}(\eta_{\gamma}||\pi)-\mathrm{KL}(\mu_{\gamma}||\pi)$ (red line) for a single sequential split step as a function of step size $\gamma$.
The KL is calculated exactly using the formula for Gaussians (B.24) and mean and covariance using the equations above. We consider the case where the target is more diffuse than the initial distribution (left plot, $C_{\pi}=100,m_{\pi}=20,C_{0}=1,m_{0}=0$) and the opposite case (right plot, $C_{\pi}=1,m_{\pi}=20,C_{0}=100,m_{0}=0$). Figure 3.1 shows that the improvement in KL due to splitting for larger $\gamma$ depends on the both the order of the operators and the on the relationship between the target and initial distribution. Specifically, in the univariate Gaussian case, when the target variance is larger than the initial variance, the ordering W-FR leads to a speed-up over the exact dynamics for all $\gamma>0$ see Figure 3.1 left panel. As expected for $\gamma$ large enough, the difference approaches zero as both the W and FR flows individually have been evaluated for a sufficiently long time so that the resulting densities are close to $\pi$. Conversely, when the target variance is smaller than the initial variance the ordering FR-W is superior, see Figure 3.1 right panel.
These numerical results are somewhat intuitive; if $\pi$ is more diffuse than $\mu_{0}$, applying the W operator first allows to increase variance and transform the initial distribution into a more diffuse one; if instead $\pi$ is more concentrated than $\mu_{0}$ applying the FR operator first allows to shrink the variance and returns a more concentrated distribution since the ratio $\pi/\mu_{0}$ will be small over most of the support. We will make this intuition more precise in the remainder of this section and extend to the multivariate Gaussian case, where we show that the difference in initial and target means plays a role also.
@@
As a first step, we present an equality for the continuous time KL decay for the WFR PDE (2.1), which to the best of our knowledge is missing from the literature. Existing KL bounds for more general $\pi$ developed in the literature are not as tight as discussed in Section 5, see e.g. Figure 5.1(left).
The exact covariance and mean evolution in (3.4)-(3.3) for the W-FR split and (3.6)-(3.5) for the FR-W split can be used in a recursive fashion to characterise the decay of $\mathrm{KL}$ distance to $\pi$ of the split schemes compared to the exact WFR flow. We will make use of the following quantities throughout,
$E_{n}:=C_{n\gamma}-C_{\pi};\quad$ $E_{n}^{\beta}:=Q_{n\gamma}-C_{\pi};\quad E_{n}^{\alpha}:=P_{n\gamma}-C_{\pi};$ (3.7)
$\varepsilon_{n}:=m_{n\gamma}-m_{\pi};\quad$ $\varepsilon_{n}^{\beta}:=b_{n\gamma}-m_{\pi};\quad\varepsilon_{n}^{\alpha}:=a_{n\gamma}-m_{\pi}.$ (3.8)
Consider Figure 3.2 which shows the ratio of Split $\mathrm{KL}$ to exact WFR $\mathrm{KL}$ for a 10D Gaussian target (see Appendix B [us_neurips] for details) with $\gamma=0.7$. For this particular case, the W-FR scheme offers a speed-up over the exact dynamics, where $\mathrm{KL}$ from the split scheme is $\approx 60\%$ of $\mathrm{KL}$ in the exact case for large $t$. Conversely, a FR-W scheme leads to a reduction in convergence speed (see Figure 3.2 (right)). Importantly, while the ratio stabilises at approximately $1.8$, both $\mathrm{KL}(\eta_{n}||\pi)$ and $\mathrm{KL}(\mu_{n\gamma}||\pi)$ approach zero so that there is no asymptotic bias.
@@
KL decay, multivariate Gaussian. Consider the multivariate Gaussian setting, i.e. $\mu_{0}(x)=\mathcal{N}(x;m_{0},C_{0})$ and $\pi(x)=\mathcal{N}(x;m_{\pi},C_{\pi})$ with $C_{0},C_{\pi}$ strictly positive definite and $C_{0}\neq C_{\pi}$. Recall $\mu_{n\gamma}$ is the solution of (2.1) at time $t=n\gamma$ and $\nu_{n}$ is the solution of the W-FR split scheme at step $n$ with step size $\gamma$. Assume the exact WFR (2.1), W-FR split scheme (2.6) and FR-W split scheme (2.7) are initialised with the same $C_{0},m_{0}$ (i.e. $E_{0},\varepsilon_{0}$ fixed across methods), so that $\mu_{0}=\nu_{0}=\eta_{0}$. In particular, assume that $E_{0}$ is either strictly negative definite or strictly positive definite. Recall $\Gamma:=C_{\pi}^{-1}+\frac{1}{2}I$. Define also the functional
@@
where $A$ denotes any $d\times d$ symmetric matrix and the matrix operator
$J_{n}(B)=e^{-n\gamma\Gamma}\left(E_{0}^{-1}+B\,C_{\pi}^{-1}(I-e^{-2n\gamma\Gamma})\right)^{-1}e^{-n\gamma\Gamma},$
for any $B\in\mathbb{R}^{d\times d}$. Then the following holds for all $n=1,2,3,\dots$,
  1. 1.
@@
    $\mathrm{KL}(\mu_{n\gamma}||\pi)$ $=\phi_{n}(E_{n}),$ (3.9)
    $E_{n}$ $=J_{n}(\Omega),$ (3.10)
    $\Omega$ $:=\frac{1}{2}\Gamma^{-1}.$ (3.11)
  2. 2.
    The recursive solution of the W-FR split PDE in Proposition 2.2 after $n$ time steps of size $\gamma$ satisfies
    $\mathrm{KL}(\nu_{n}||\pi)$ $=\phi_{n}(E_{n}^{\beta}),$ (3.12)
    $E_{n}^{\beta}$ $=J_{n}(\Omega^{\beta}),$ (3.13)
    $\Omega^{\beta}$ $:=(1-e^{\gamma})(I-e^{2\gamma\Gamma})^{-1}.$ (3.14)
  3. 3.
    The recursive solution of the FR-W split PDE in Proposition 2.3 after $n$ time steps of size $\gamma$ satisfies
    $\mathrm{KL}(\eta_{n\gamma}||\pi)$ $=\phi_{n}(E_{n}^{\alpha}),$ (3.15)
    $E_{n}^{\alpha}$ $=J_{n}(\Omega^{\alpha}),$ (3.16)
Lemma 3.1 shows that the change in KL decay due to splitting is, as discussed previously, entirely due to differences in covariance estimation. Despite the considerable difference in structure of $C_{\gamma}$ vs $Q_{\gamma}$ or $P_{\gamma}$ (c.f. (3.1) with $t=\gamma$ and (3.3), (3.5)), the $n$-step iterations lead to expressions that more clearly highlight similarities between the exact and sequential splitting. More specifically, notice that the difference quantities $E_{n}$ and $E_{n}^{\beta},E_{n}^{\alpha}$ are identical up to the $\Omega$ and $\Omega^{\beta},\Omega^{\alpha}$ factors.
@@
    Suppose $E_{0}\succ 0$. Then $E_{n}\succ 0,E_{n}^{\beta}\succ 0$ and $E_{n}^{\alpha}\succ 0$ for all $n=1,2,3\dots$.
@@
    Suppose $E_{0}^{-1}C_{\pi}\prec-\Omega$, $E_{0}^{-1}C_{\pi}\prec-\Omega^{\beta}$ and $E_{0}^{-1}C_{\pi}\prec-\Omega^{\alpha}$. Then $E_{n}\prec 0,E_{n}^{\beta}\prec 0$ and $E_{n}^{\alpha}\prec 0$ for all $n=1,2,3,\dots$.

### v2

@@ 2.2 Variational Formula: FR-W
The reverse scheme described in (2.7) requires a more involved analysis than in Proposition 2.2, due to the lack of invertibility of $S_{\textrm{W}}$ in general. We instead rely on techniques used in developing Alekseev-Groebner type formulae, in particular, the formal Lie calculus, as is common in the analysis of operator splitting schemes [DT13, HV03]. We rely on the shorthand notation $\eta_{\gamma}$ to refer to $\eta_{1}(x;\gamma)$. We have the next proposition for the FR-W scheme. Its proof can be found in Appendix A.2.
@@
Sequential split PDE: FR-W. Let us denote $\eta_{1}=\eta_{1}(x;\gamma)$ and let Assumption 1 hold. The variation of one sequential split step of size $\gamma$ in the order FR-W (2.7) corresponds to the PDE
$\partial_{\gamma}\eta_{1}=f_{\textrm{W}}(\eta_{1})+f_{\textrm{FR}}(\eta_{1})+\int_{0}^{\gamma}S_{\textrm{W}}\left(\gamma-\tau,-f_{\textrm{P}}(\eta_{\tau})\right)d\tau,$ (2.11)
@@
$\partial_{\gamma}\eta_{1}=f_{\textrm{W}}(\eta_{1})+f_{\textrm{FR}}(\eta_{1})+\sum_{k=1}^{\infty}\frac{\gamma^{k}}{k!}\eta_{1}\left(g_{k}(\eta_{1})-\mathbb{E}_{\eta_{1}}[g_{k}]\right),$ (2.12)
@@
$g_{0}(\eta_{1})$ $=-g(\eta_{1})=-\log\frac{\eta_{1}}{\pi},$
$g_{k}(\eta_{1})$ $=-g_{k-1}^{\prime}f_{\textrm{W}}(\eta_{1})+\nabla\log\frac{\eta_{1}}{\pi}\cdot\nabla g_{k-1}(\eta_{1})+\frac{1}{\eta_{1}}\nabla\cdot(\eta_{1}\nabla g_{k-1}(\eta_{1})),\quad\text{for}\enskip k=2,3,\dots$
and $g_{k-1}^{\prime}f_{\textrm{W}}(\eta_{1})$ denotes the directional derivative of $g_{k-1}$ at $\eta_{1}$ in the direction $f_{\textrm{W}}(\eta_{1})$.
The expression in (2.12) is obtained by writing the perturbation term in (2.11) using the Lie commutator in (1.5) and higher order commutators. We work with this formulation as the inverse of the solution operator $S_{\textrm{W}}$ does not lead to a tractable form, when it exists. A similar approach can be used to obtain the perturbation term in Proposition 2.2 with $g_{k}=|\nabla g|^{2}$ for all $k=1,2,3,\dots$. We do not use this approach in the proof of Proposition 2.2 but exploit the invertibility of the unnormalised FR solution operator.
As Proposition 2.3 shows, the evolution of the splitting solution obtained with FR-W does not correspond to that obtained in Proposition 2.2 for W-FR. In the multivariate Gaussian setting, i.e. with $\mu_{0}(x)=\mathcal{N}(x;m_{0},C_{0})$ and $\pi(x)=\mathcal{N}(x;m_{\pi},C_{\pi})$, a more interpretable PDE reminiscent of (2.9) can be obtained by direct evaluation of (2.12) as
$\partial_{\gamma}\eta_{1}$ $=f_{\textrm{W}}(\eta_{1})+f_{\textrm{FR}}(\eta_{1})-\eta_{1}\left(g_{\text{FRW}}(\eta_{1})-\mathbb{E}_{\eta_{1}}[g_{\text{FRW}}(\eta_{1})]\right),$ (2.13)
@@ 3 Splitting & convergence speed: multivariate Gaussian case
where $\Gamma:=C_{\pi}^{-1}+\frac{1}{2}I$. A derivation of (3.1)-(3.2) can be found in Lemma B.2 (letting $M_{t}=0$), which extends the derivation of $C_{t}^{-1}$ provided in [LMT+25]. To the best of our knowledge, the analytic solution for $m_{t}$ in (3.2) has not yet appeared in the literature.
Similarly, the splitting schemes yield Gaussian solutions at each time, as it is well-known that both W and FR flows preserve Gaussianity. We denote by $b_{n},Q_{n}$ the mean and covariance respectively at step $n$ of the W-FR split scheme with step size $\gamma$ and $a_{n},P_{n}$ the mean and covariance at step $n$ of the FR-W split.
The similarity in structure of the sequential split PDEs (2.9) and (2.13) in the Gaussian setting also allows for obtaining analytic solutions of the mean and covariance in a similar fashion. An application of Lemma B.2 with $\gamma$ in place of $t$ and setting $M_{\gamma}=(e^{\gamma}-1)I$, gives
$Q_{1}$ $=C_{\pi}+(e^{\Gamma\gamma}(C_{0}-C_{\pi})^{-1}e^{\Gamma\gamma}+(e^{\gamma}-1)C_{\pi}^{-1})^{-1},$ (3.3)
$b_{1}$ $=m_{\pi}+(Q_{\gamma}-C_{\pi})e^{\gamma C_{\pi}^{-1}}(C_{0}-C_{\pi})^{-1}(m_{0}-m_{\pi}).$ (3.4)
for the mean $b_{1}$ and covariance $Q_{1}$ of a single step of W-FR of size $\gamma$. Similarly, setting $M_{\gamma}=-\frac{1}{2}C_{\pi}(e^{2\gamma C_{\pi}^{-1}}-I)$ in Lemma B.2 yields
$P_{1}$ $=C_{\pi}+\left(e^{\Gamma\gamma}(C_{0}-C_{\pi})^{-1}e^{\Gamma\gamma}+(e^{\gamma}-1)C_{\pi}^{-1}e^{2\gamma C_{\pi}^{-1}}\right)^{-1}$ (3.5)
$a_{1}$ $=m_{\pi}+(P_{\gamma}-C_{\pi})e^{\gamma C_{\pi}^{-1}}(C_{0}-C_{\pi})^{-1}(m_{0}-m_{\pi}).$ (3.6)
for the mean $a_{1}$ and covariance $P_{1}$ of a single step of FR-W of size $\gamma$.
The structure of the mean for both splitting schemes is exactly the same as that obtained for the exact WFR flow with $C_{\gamma}$ replaced by $Q_{1},P_{1}$ respectively. This demonstrates that in the Gaussian case the speed-up due to splitting is driven primarily through the improvement in covariance estimation. In particular, notice the difference in structure between $C_{\gamma}$ vs $Q_{1}$ and $P_{1}$. The difference due to operator ordering is more subtle, as $Q_{1}$ and $P_{1}$ differ only by a single factor $e^{2\gamma C_{\pi}^{-1}}$. As will be shown later in this section, this factor does not necessarily lead to acceleration in all cases, and depends on the relationship between $\mu_{0}$ and $\pi$.
@@
Before presenting the main analytic results, consider Figure 3.1 which compares KL distance from $\pi$ after a single sequential split step and the same for the exact WFR in the univariate Gaussian setting. More specifically, it shows $\mathrm{KL}(\nu_{1}||\pi)-\mathrm{KL}(\mu_{\gamma}||\pi)$ (blue line) and $\mathrm{KL}(\eta_{1}||\pi)-\mathrm{KL}(\mu_{\gamma}||\pi)$ (red line) for a single sequential split step as a function of step size $\gamma$. The KL is calculated exactly using the formula for Gaussians (2.16) and mean and covariance using the equations above. We consider the case where the target is more diffuse than the initial distribution (left plot, $C_{\pi}=100,m_{\pi}=20,C_{0}=1,m_{0}=0$) and the opposite case (right plot, $C_{\pi}=1,m_{\pi}=20,C_{0}=100,m_{0}=0$). Figure 3.1 shows that the improvement in KL due to splitting for larger $\gamma$ depends on the both the order of the operators and the on the relationship between the target and initial distribution. Specifically, in the univariate Gaussian case, when the target variance is larger than the initial variance, the ordering W-FR leads to a speed-up over the exact dynamics for all $\gamma>0$, see Figure 3.1 left panel. As expected for $\gamma$ large enough, the difference approaches zero as both the W and FR flows individually have been evaluated for a sufficiently long time so that the resulting densities are close to $\pi$. Conversely, when the target variance is smaller than the initial variance the ordering FR-W is superior, see Figure 3.1 right panel.
These numerical results are somewhat intuitive; if $\pi$ is more diffuse than $\mu_{0}$, applying the W operator first allows to increase variance and transform the initial distribution into a more diffuse one; if instead $\pi$ is more concentrated than $\mu_{0}$ applying the FR operator first allows to shrink the variance and returns a more concentrated distribution since the ratio $\pi/\mu_{0}$ will be small over most of the support. We will make this intuition more precise in the remainder of this section and extend to the multivariate setting, where we show that the difference in initial and target means plays a role also.
@@
As a first step, we present an equality for the continuous time KL decay for the WFR PDE (2.1), which to the best of our knowledge is missing from the literature. Existing KL bounds for more general $\pi$ developed in the literature are not as tight as discussed in Section 5.1, see e.g. Figure 5.1(left). The exact covariance and mean evolution in (3.4)-(3.3) for the W-FR split and (3.6)-(3.5) for the FR-W split can be used in a recursive fashion to characterise the decay of $\mathrm{KL}$ distance to $\pi$ of the split schemes compared to the exact WFR flow. We will make use of the following quantities throughout,
$E_{n}:=C_{n\gamma}-C_{\pi};\quad$ $E_{n}^{\beta}:=Q_{n}-C_{\pi};\quad E_{n}^{\alpha}:=P_{n}-C_{\pi};$ (3.7)
$\varepsilon_{n}:=m_{n\gamma}-m_{\pi};\quad$ $\varepsilon_{n}^{\beta}:=b_{n}-m_{\pi};\quad\varepsilon_{n}^{\alpha}:=a_{n}-m_{\pi}.$ (3.8)
Consider Figure 3.2 which shows the ratio of split $\mathrm{KL}$ to exact WFR $\mathrm{KL}$ for a 10D Gaussian target with $\gamma=0.7$. For this particular case, the W-FR scheme offers a speed-up over the exact dynamics, where $\mathrm{KL}$ from the split scheme is $\approx 60\%$ of $\mathrm{KL}$ in the exact case for large $t$. Conversely, a FR-W scheme leads to a reduction in convergence speed (see Figure 3.2 (right)). Importantly, while the ratio stabilises at approximately $1.8$, both $\mathrm{KL}(\eta_{n}||\pi)$ and $\mathrm{KL}(\mu_{n\gamma}||\pi)$ approach zero so that there is no time asymptotic bias.
@@
KL decay, multivariate Gaussian. Consider the multivariate Gaussian setting, i.e. $\mu_{0}(x)=\mathcal{N}(x;m_{0},C_{0})$ and $\pi(x)=\mathcal{N}(x;m_{\pi},C_{\pi})$ with $C_{0},C_{\pi}$ strictly positive definite and $C_{0}\neq C_{\pi}$. Assume the exact WFR (2.1), W-FR split scheme (2.6) and FR-W split scheme (2.7) are initialised with the same $C_{0},m_{0}$ (i.e. $E_{0},\varepsilon_{0}$ fixed across methods), so that $\mu_{0}=\nu_{0}=\eta_{0}$. Recall $\Gamma:=C_{\pi}^{-1}+\frac{1}{2}I$. Define also
@@
$J_{n}(B)$ $:=e^{-n\gamma\Gamma}\left(E_{0}^{-1}+B\,C_{\pi}^{-1}(I-e^{-2n\gamma\Gamma})\right)^{-1}e^{-n\gamma\Gamma},$
where $A$ denotes any $d\times d$ symmetric matrix and $B\in\mathbb{R}^{d\times d}$. Finally, let
$\Omega$ $:=\frac{1}{2}\Gamma^{-1}$ (3.9)
$\Omega^{\beta}$ $:=(1-e^{\gamma})(I-e^{2\gamma\Gamma})^{-1},$ (3.10)
@@
$\mathrm{KL}(\mu_{n\gamma}||\pi)=\phi_{n}(E_{n}),$ $\enskip E_{n}=J_{n}(\Omega)$
Similarly, for the split schemes, $\mathrm{KL}(\nu_{n}||\pi)=\phi_{n}(E_{n}^{\beta}),\enskip E_{n}^{\beta}=J_{n}(\Omega^{\beta}),$ and $\mathrm{KL}(\eta_{n}||\pi)=\phi_{n}(E_{n}^{\beta}),\enskip E_{n}^{\alpha}=J_{n}(\Omega^{\alpha})$.
Lemma 3.1 shows that the change in KL decay due to splitting is, as discussed previously, entirely due to differences in covariance estimation. Despite the considerable difference in structure of $C_{\gamma}$ vs $Q_{\gamma}$ or $P_{\gamma}$ (c.f. (3.1) with $t=\gamma$ and (3.3), (3.5)), the $n$-step iterations lead to expressions that more clearly highlight similarities between the exact and sequential splitting. More specifically, notice that the difference quantities $E_{n}$ and $E_{n}^{\beta},E_{n}^{\alpha}$ are identical up to factors depending on $\Omega$ and $\Omega^{\beta},\Omega^{\alpha}$.
@@
[…]
@@
[…]